# Position: Untrained Machine Learning for Anomaly Detection

### v2

@@ 1 Introduction
Anomaly detection is a crucial research problem within the machine learning and quality control communities. Recently, there has been significant research focused on using machine learning for anomaly detection. Based on the availability of labels, anomaly detection methods can be categorized into three types: unsupervised ([Zhang et al.(2024)Zhang, Li, Wu, Li, Lin, Hu, Li, and Jiang]), supervised ([Jiang et al.(2023)Jiang, Hou, Zheng, Han, Huang, Wen, Hu, and Zhao]), and semi-supervised anomaly detection ([Qiao et al.(2024)Qiao, Wen, Li, Lim, and Pang]). Unsupervised anomaly detection is particularly popular because it does not require labeled data, which is often scarce, allowing the algorithm to independently identify anomalies without the need for predefined labels. In contrast, supervised anomaly detection relies on labeled data, which limits the algorithm to detect only those anomalies that were encountered during training. Semi-supervised anomaly detection combines elements of both supervised and unsupervised approaches, enabling the handling of some labeled data alongside large amounts of unlabeled data. Among these three categories, unsupervised anomaly detection is particularly valuable because it eliminates the need for costly data labeling. Here in this paper, anomaly mainly refers as surface anomaly.
@@
Notably, untrained methods diverge fundamentally from the popular zero-shot and few-shot learning paradigms. While both untrained and zero-shot methods operate without task-specific training data, their fundamental approaches differ significantly. Zero-shot learning leverages pre-trained models on auxiliary datasets through semantic mappings to recognize unseen categories, making it suitable when prior defect templates are available but real anomaly samples are absent. Few-shot learning achieves rapid adaptation using minimal labeled samples, though its performance is strongly influenced by sample quality and quantity. In contrast, untrained methods bypass all training data entirely, achieving anomaly detection solely through geometric priors (e.g., surface smoothness, low-rankness) and physical constraints (e.g., sparsity assumptions). This distinction is particularly critical in personalized manufacturing contexts, where three critical constraints exist: (1) uniqueness of products eliminates the feasibility of obtaining labeled training data; (2) heterogeneous 3D scanners with varying point cloud densities preclude the use of pre-trained models; and (3) absence of CAD models for most customized designs invalidates template-based inspection methods. By directly analyzing geometric structural features from single samples and balancing computational efficiency with zero reliance on historical data, untrained methods maintain detection efficacy under extreme data scarcity, offering an innovative solution to industrial challenges of ”zero sample availability” in personalized manufacturing and healthcare scenarios.
@@
To address these challenges, several initial efforts have been made, including local geometry-based methods [Wang et al.(2023b)Wang, Sun, Jin, Kong, and Yue],decomposition-based approaches [Tao & Du(2025)Tao and Du], and statistical latent variable inference methods [Tao et al.(2023a)Tao, Du, and Chang]. A more detailed review will be provided in the following section. However, many research problems still need to be addressed. To advance the current state of untrained machine learning for anomaly detection, this paper first formally defines the anomaly detection problem via untrained machine learning, and three potential research methodology frameworks are proposed. Our contribution is illustrated in Figure 1. Three specific examples for each framework are also provided. Numerical studies are conducted to demonstrate the effectiveness and computational efficiency of untrained methods compared to existing training-based unsupervised learning methods. Finally, the paper concludes with a discussion and future outlook for untrained machine learning anomaly detection.
@@ 2 Paper Review
In the realm of manufacturing, anomaly detection using 3D point cloud data has witnessed a significant shift towards unsupervised methods, primarily due to the scarcity of annotated data. These unsupervised approaches can be broadly categorized into training-based methods and untrained methods, each with its own characteristics and applications. Training-based unsupervised methods operate under the assumption that sufficient anomaly-free training samples are available. This enables them to learn the intrinsic patterns of normal surfaces, which are then used to detect anomalies. There are two main techniques within this category: feature embedding and reconstruction. Examples can be referred to [Cao et al.(2024b)Cao, Xu, and Shen, Chu et al.(2023)Chu, Liu, Hsieh, Chen, and Liu, Horwitz & Hoshen(2023b)Horwitz and Hoshen, Wang et al.(2023a)Wang, Peng, Zhang, Yi, Wang, and Wang]. Feature embedding-based methods involve a two-step process. First, latent features are extracted from anomaly-free training data. For example, in memory bank-based methods like Back To the Feature (BTF) [Horwitz & Hoshen(2023b)Horwitz and Hoshen], traditional descriptors such as FPFH are used to extract features from training point cloud patches, and these features are stored in a memory bank. The memory bank is then downsampled to represent the distribution of normal features. During inference, features that deviate significantly from this distribution are flagged as anomalies. Another approach is to use knowledge distillation (KD). Methods like 3D-ST [Bergmann & Sattlegger(2023)Bergmann and Sattlegger] transfer knowledge from a pre-trained teacher network (e.g., RandLA-Net [Hu et al.(2022)Hu, Yang, Xie, Rosa, Guo, Wang, Trigoni, and Markham]) to a student network. The student network is trained on anomaly-free data to mimic the teacher’s output, and discrepancies between the two are used to calculate anomaly scores during inference.
@@
Reconstruction-based methods aim to achieve anomaly detection at the point level. Autoencoder-based methods, such as EasyNet [Chen et al.(2023)Chen, Xie, Liu, Wang, Luo, Wang, and Zheng], use a feature encoder and decoder to reconstruct 3D point cloud data. Trained only on anomaly-free data, they calculate anomaly scores based on the discrepancies between the input and the reconstructed data. For instance, EasyNet employs a multi-scale, multi-modality feature encoder-decoder for 3D depth map reconstruction, enabling real-time detection. However, some autoencoder-based methods like EasyNet and Cheating Depth [Zavrtanik et al.(2024)Zavrtanik, Kristan, and Skoaj] require structured depth maps as input and struggle with unstructured point clouds. To address this, [Li et al.(2023)Li, Xu, Gu, Zheng, Gao, and Wu] proposed the self-supervised Iterative Mask Reconstruction Network (IMRNet) for unstructured point cloud reconstruction and anomaly detection. Principal component analysis (PCA)-based methods, like those by [Von Enzberg & Al-Hamadi(2016)Von Enzberg and Al-Hamadi] and [Zhang et al.(2018)Zhang, Zou, Lin, Xu, He, Gui, and Li], identify normal patterns through principal components from training data and use them to reconstruct test samples for anomaly detection. Despite their effectiveness, training-based methods are limited by the requirement for extensive anomaly-free training data. They also face challenges in accurately detecting anomalies with sparse properties or similar patterns to normal data. Feature embedding-based methods are good at localizing anomalies but have difficulty in obtaining accurate anomaly boundaries, while reconstruction-based methods can get more accurate anomaly boundaries but struggle with complex surface reconstruction, leading to higher false positive rates [Masuda et al.(2023)Masuda, Hachiuma, Fujii, Saito, and Sekikawa, Zavrtanik et al.(2024)Zavrtanik, Kristan, and Skoaj, Li et al.(2023)Li, Xu, Gu, Zheng, Gao, and Wu, Roth et al.(2022)Roth, Pemula, Zepeda, Scholkopf, Brox, and Gehler].
@@
Untrained methods, on the other hand, do not require training on a large dataset. They rely on prior knowledge to model the normal surface or possible anomalies ([Jovanevi et al.(2017)Jovanevi, Pham, Orteu, Gilblas, Harvent, Maurice, and Brthes]). Local geometry-based methods explore local geometric characteristics to detect anomalies. For example, [Jovanevi et al.(2017)Jovanevi, Pham, Orteu, Gilblas, Harvent, Maurice, and Brthes] used a region-growing segmentation algorithm with local normal and curvature data to segment point cloud airplane surfaces. [Wei et al.(2021)Wei, Chengyao, Shilin, and Kailin] developed local features to compute anomaly scores, and [Miao et al.(2022)Miao, Fu, Wu, Hao, Li, Hao, and Zhou] used FPFH and normal vector aggregation for defect detection on gas turbine blades. Global geometry-based methods utilize the global shapes of manufacturing parts. Statistical-based methods, such as the one proposed by [Tao et al.(2023a)Tao, Du, and Chang], make assumptions about the shape of the product, like low-rankness and smoothness, and formulate the anomaly detection problem within a probabilistic framework. CAD model-based methods compare the point cloud with CAD models through rigid registration. [Zhao et al.(2023)Zhao, Li, Xiao, and He] used the standard iterative closest point (ICP) algorithm for defect detection of 3D printing products, and various improvements have been made to the ICP algorithm, such as the octree-based registration algorithm by [He et al.(2023)He, Ma, Li, Hao, Wang, and Wang] to accelerate the process. Untrained methods can handle one single sample directly, which is an advantage in scenarios with limited data.
@@ 3.1 Latent Variable Inference Framework
Figure 2: Example of Latent Variable Inference Framework [Tao et al.(2023b)Tao, Du, and Chang]
An example of the Latent Variable Inference framework is provided by [Tao et al.(2023a)Tao, Du, and Chang], whom establishes probabilistic relationships between point locations and neighborhood smoothness, as shown in Figure 2. This approach models the dependencies between surface points and their local geometric features, allowing for more precise anomaly identification in complex 3D structures. More details can be found in Appendix LABEL:apLV.
@@ 3.2 Decomposition Framework
Figure 3: Example of Decomposition Framework [Tao & Du(2025)Tao and Du]
An example of the Decomposition-based framework is provided by [Tao & Du(2025)Tao and Du], which achieves sparse anomaly recovery through graph-based smoothness constraints and group LOG penalties. This method effectively isolates anomalous components in the 3D point cloud data as shown in Figure 3. The framework offers a robust strategy for identifying subtle surface deviations within large datasets. More details can be referred to the Appendix LABEL:ApD.
@@ 3.3 Local Geometry Framework
There are various local-geometry based feature learning methods mentioned in [Du et al.(2025)Du, Tao, Cao, and Tsung], and key ideas are utilizing local shape descriptors to extract the geometry feature from the neighbourhood. More details can be referred to [Du et al.(2025)Du, Tao, Cao, and Tsung]. Due to the local modeling of the object surfaces, local-geometry-based methods usually cannot detect anomaly boundaries very accurately, so the authors suggest prioritizing the first two frameworks. However, when the assumptions of the latent variable inference framework and decomposition framework fail, the local-geometry-based framework can serve as an alternative, which is more general in real applications.
Figure 4: Example of Local-Geometry based Framework, the influence region diagram for a Point Feature Histogram. The query point (orange) and its k-neighbors (yellow) are fully interconnected in a dash circle. [Rusu et al.(2009-05)Rusu, Blodow, and Beetz]
An example of the Local Geometry Framework framework is provided by [Rusu et al.(2009)Rusu, Blodow, and Beetz], which exemplifies local geometry analysis through Darboux frame construction and feature histogram computation, as ahown in Figure 4. By analyzing point neighborhoods using invariant geometric features, the method provides a powerful tool for local anomaly detection in 3D point clouds. More details can be found in Appendix LABEL:apL.
@@ 4 Experiments and Analysis
  • Training-based Methods: Compared with memory bank approaches utilizing pre-trained features, specifically BTF ([Horwitz & Hoshen(2023a)Horwitz and Hoshen]) and CPMF ([Cao et al.(2024a)Cao, Xu, and Shen]), the memory bank was constructed using 300 training samples with corresponding construction time shown in Table 2.
Evaluation Metrics The evaluation metrics (pointwise) are defined as: $\text{FOR}=\frac{\text{FN}}{\text{FN}+\text{TN}}$, $\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}}$, $\text{BA}=\frac{1}{2}(\frac{\text{TP}}{\text{TP}+\text{FN}}+\frac{\text{TN}}{\text{TN}+\text{FP}})$, and $\text{DICE}=\frac{2\text{TP}}{2\text{TP}+\text{FP}+\text{FN}}$, where TP, TN, FP, and FN represent true positives, true negatives, false positives, and false negatives, and positive refers to a data point detected as anomalous. These evaluation indices are commonly used indices for anomaly detection by using 3D point cloud data ([Du et al.(2025)Du, Tao, Cao, and Tsung]).

### v3

@@ 1 Introduction
Anomaly detection is a crucial research problem within the machine learning and quality control communities. Recently, there has been significant research focused on using machine learning for anomaly detection. Based on the availability of labels, anomaly detection methods can be categorized into three types: unsupervised (Zhang et al. (2024)), supervised (Jiang et al. (2023)), and semi-supervised anomaly detection (Qiao et al. (2024)). Unsupervised anomaly detection is particularly popular because it does not require labeled data, which is often scarce, allowing the algorithm to independently identify anomalies without the need for predefined labels. In contrast, supervised anomaly detection relies on labeled data, which limits the algorithm to detect only those anomalies that were encountered during training. Semi-supervised anomaly detection combines elements of both supervised and unsupervised approaches, enabling the handling of some labeled data alongside large amounts of unlabeled data. Among these three categories, unsupervised anomaly detection is particularly valuable because it eliminates the need for costly data labeling. Here in this paper, anomaly mainly refers as surface anomaly.
@@
Notably, untrained methods diverge fundamentally from the popular zero-shot and few-shot learning paradigms. While both untrained and zero-shot methods operate without task-specific training data, their fundamental approaches differ significantly. Zero-shot learning leverages pre-trained models on auxiliary datasets through semantic mappings to recognize unseen categories, making it suitable when prior defect templates are available but real anomaly samples are absent. Few-shot learning achieves rapid adaptation using minimal labeled samples, though its performance is strongly influenced by sample quality and quantity. In contrast, untrained methods bypass all training data entirely, achieving anomaly detection solely through geometric priors (e.g., surface smoothness, low-rankness) and physical constraints (e.g., sparsity assumptions). This distinction is particularly critical in personalized manufacturing contexts, where three critical constraints exist: (1) uniqueness of products eliminates the feasibility of obtaining labeled training data; (2) heterogeneous 3D scanners with varying point cloud densities preclude the use of pre-trained models; and (3) absence of CAD models for most customized designs invalidates template-based inspection methods. By directly analyzing geometric structural features from single samples and balancing computational efficiency with zero reliance on historical data, untrained methods maintain detection efficacy under extreme data scarcity, offering an innovative solution to industrial challenges of "zero sample availability" in personalized manufacturing and healthcare scenarios.
@@
To address these challenges, several initial efforts have been made, including local geometry-based methods Wang et al. (2023a),decomposition-based approaches Tao and Du (2025), and statistical latent variable inference methods Tao et al. (2023a). A more detailed review will be provided in the following section. However, many research problems still need to be addressed. To advance the current state of untrained machine learning for anomaly detection, this paper first formally defines the anomaly detection problem via untrained machine learning, and three potential research methodology frameworks are proposed. Our contribution is illustrated in Figure 1. Three specific examples for each framework are also provided. Numerical studies are conducted to demonstrate the effectiveness and computational efficiency of untrained methods compared to existing training-based unsupervised learning methods. Finally, the paper concludes with a discussion and future outlook for untrained machine learning anomaly detection.
@@ 2 Paper Review
In the realm of manufacturing, anomaly detection using 3D point cloud data has witnessed a significant shift towards unsupervised methods, primarily due to the scarcity of annotated data. These unsupervised approaches can be broadly categorized into training-based methods and untrained methods, each with its own characteristics and applications. Training-based unsupervised methods operate under the assumption that sufficient anomaly-free training samples are available. This enables them to learn the intrinsic patterns of normal surfaces, which are then used to detect anomalies. There are two main techniques within this category: feature embedding and reconstruction. Examples can be referred to Cao et al. (2024a); Chu et al. (2023); Horwitz and Hoshen (2023a); Wang et al. (2023b). Feature embedding-based methods involve a two-step process. First, latent features are extracted from anomaly-free training data. For example, in memory bank-based methods like Back To the Feature (BTF) Horwitz and Hoshen (2023a), traditional descriptors such as FPFH are used to extract features from training point cloud patches, and these features are stored in a memory bank. The memory bank is then downsampled to represent the distribution of normal features. During inference, features that deviate significantly from this distribution are flagged as anomalies. Another approach is to use knowledge distillation (KD). Methods like 3D-ST Bergmann and Sattlegger (2023) transfer knowledge from a pre-trained teacher network (e.g., RandLA-Net Hu et al. (2022)) to a student network. The student network is trained on anomaly-free data to mimic the teacher’s output, and discrepancies between the two are used to calculate anomaly scores during inference.
@@
Reconstruction-based methods aim to achieve anomaly detection at the point level. Autoencoder-based methods, such as EasyNet Chen et al. (2023), use a feature encoder and decoder to reconstruct 3D point cloud data. Trained only on anomaly-free data, they calculate anomaly scores based on the discrepancies between the input and the reconstructed data. For instance, EasyNet employs a multi-scale, multi-modality feature encoder-decoder for 3D depth map reconstruction, enabling real-time detection. However, some autoencoder-based methods like EasyNet and Cheating Depth Zavrtanik et al. (2024) require structured depth maps as input and struggle with unstructured point clouds. To address this, Li et al. (2023) proposed the self-supervised Iterative Mask Reconstruction Network (IMRNet) for unstructured point cloud reconstruction and anomaly detection. Principal component analysis (PCA)-based methods, like those by Von Enzberg and Al-Hamadi (2016) and Zhang et al. (2018), identify normal patterns through principal components from training data and use them to reconstruct test samples for anomaly detection. Despite their effectiveness, training-based methods are limited by the requirement for extensive anomaly-free training data. They also face challenges in accurately detecting anomalies with sparse properties or similar patterns to normal data. Feature embedding-based methods are good at localizing anomalies but have difficulty in obtaining accurate anomaly boundaries, while reconstruction-based methods can get more accurate anomaly boundaries but struggle with complex surface reconstruction, leading to higher false positive rates Masuda et al. (2023); Zavrtanik et al. (2024); Li et al. (2023); Roth et al. (2022).
@@
Untrained methods, on the other hand, do not require training on a large dataset. They rely on prior knowledge to model the normal surface or possible anomalies (Jovanevi et al. (2017)). Local geometry-based methods explore local geometric characteristics to detect anomalies. For example, Jovanevi et al. (2017) used a region-growing segmentation algorithm with local normal and curvature data to segment point cloud airplane surfaces. Wei et al. (2021) developed local features to compute anomaly scores, and Miao et al. (2022) used FPFH and normal vector aggregation for defect detection on gas turbine blades. Global geometry-based methods utilize the global shapes of manufacturing parts. Statistical-based methods, such as the one proposed by Tao et al. (2023a), make assumptions about the shape of the product, like low-rankness and smoothness, and formulate the anomaly detection problem within a probabilistic framework. CAD model-based methods compare the point cloud with CAD models through rigid registration. Zhao et al. (2023) used the standard iterative closest point (ICP) algorithm for defect detection of 3D printing products, and various improvements have been made to the ICP algorithm, such as the octree-based registration algorithm by He et al. (2023) to accelerate the process. Untrained methods can handle one single sample directly, which is an advantage in scenarios with limited data.
@@ 3.1 Latent Variable Inference Framework
Figure 2: Example of Latent Variable Inference Framework Tao et al. (2023b)
An example of the Latent Variable Inference framework is provided by Tao et al. (2023a), whom establishes probabilistic relationships between point locations and neighborhood smoothness, as shown in Figure 2. This approach models the dependencies between surface points and their local geometric features, allowing for more precise anomaly identification in complex 3D structures. More details can be found in Appendix A.1.
@@ 3.2 Decomposition Framework
Figure 3: Example of Decomposition Framework Tao and Du (2025)
An example of the Decomposition-based framework is provided by Tao and Du (2025), which achieves sparse anomaly recovery through graph-based smoothness constraints and group LOG penalties. This method effectively isolates anomalous components in the 3D point cloud data as shown in Figure 3. The framework offers a robust strategy for identifying subtle surface deviations within large datasets. More details can be referred to the Appendix A.2.
@@ 3.3 Local Geometry Framework
There are various local-geometry based feature learning methods mentioned in Du et al. (2025), and key ideas are utilizing local shape descriptors to extract the geometry feature from the neighbourhood. More details can be referred to Du et al. (2025). Due to the local modeling of the object surfaces, local-geometry-based methods usually cannot detect anomaly boundaries very accurately, so the authors suggest prioritizing the first two frameworks. However, when the assumptions of the latent variable inference framework and decomposition framework fail, the local-geometry-based framework can serve as an alternative, which is more general in real applications.
Figure 4: Example of Local-Geometry based Framework, the influence region diagram for a Point Feature Histogram. The query point (orange) and its k-neighbors (yellow) are fully interconnected in a dash circle. Rusu et al. (2009-05)
An example of the Local Geometry Framework framework is provided by Rusu et al. (2009), which exemplifies local geometry analysis through Darboux frame construction and feature histogram computation, as ahown in Figure 4. By analyzing point neighborhoods using invariant geometric features, the method provides a powerful tool for local anomaly detection in 3D point clouds. More details can be found in Appendix A.3.
@@ 4 Experiments and Analysis
  • Training-based Methods: Compared with memory bank approaches utilizing pre-trained features, specifically BTF (Horwitz and Hoshen (2023b)) and CPMF (Cao et al. (2024b)), the memory bank was constructed using 300 training samples with corresponding construction time shown in Table 2.
Evaluation Metrics The evaluation metrics (pointwise) are defined as: $\text{FOR}=\frac{\text{FN}}{\text{FN}+\text{TN}}$, $\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}}$, $\text{BA}=\frac{1}{2}(\frac{\text{TP}}{\text{TP}+\text{FN}}+\frac{\text{TN}}{\text{TN}+\text{FP}})$, and $\text{DICE}=\frac{2\text{TP}}{2\text{TP}+\text{FP}+\text{FN}}$, where TP, TN, FP, and FN represent true positives, true negatives, false positives, and false negatives, and positive refers to a data point detected as anomalous. These evaluation indices are commonly used indices for anomaly detection by using 3D point cloud data (Du et al. (2025)).